# Noncommutative metric geometry of quantum circle bundles

### v1

@@ 10. The quantum spheres and the quantum projective spaces
We are also interested in the $q$-deformed version of the special unitary group $C(SU_{q}(r+1))$ which we refer to as the $C^{*}$-algebraic quantum special unitary group. This compact quantum group was introduced by Stanisław Woronowicz in [Wor87b, Wor87a, Wor88] and it contains the $C^{*}$-algebraic quantum sphere $C(S_{q}^{2r+1})$ as a unital $C^{*}$-subalgebra. Using the description of $C(SU_{q}(r+1))$ from [KlSc97, Chapter 9.2] the relevant map is given by $z_{i}\mapsto u_{r+1,i}$ for all $i\in\{1,\ldots,r+1\}$. The injectivity of this unital *-homomorphism is described in [VaSo90, She97, HoSz02] and a detailed argument can also be found in the review [D’A23, Section 3.3] following the scheme of [MiKa23].
@@ 10.2. Noncommutative metrics on the quantum spheres
We are now going to show that Assumption 7.5 is satisfied in the context of the higher Vaksman-Soibelman quantum spheres. This leads us straight to the last main result of this paper where we show that each of the quantum spheres can be endowed with the structure of a compact quantum metric space arising from noncommutative geometric data. In particular, we generalize one of the main results from [KaKy22] regarding quantum metrics on quantum $SU(2)$. More precisely, our methods yield a new proof of [KaKy22, Theorem 5.23] in the special case where the parameter $t$ (appearing in the statement of [KaKy22, Theorem 5.23]) is equal to one.
Our starting point is the $C^{*}$-algebraic quantum sphere $A:=C(S_{q}^{2r+1})$ which is equipped with the strongly continuous circle action $\sigma$ satisfying that $\sigma_{\lambda}(z_{i})=\lambda z_{i}$ for all $i\in\{1,\ldots,r+1\}$. We are moreover considering the coordinate algebra for the quantum sphere, $\mathcal{A}:=\mathcal{O}(S_{q}^{2r+1})$, which is a norm-dense unital *-subalgebra of $C(S_{q}^{2r+1})$. On top of this data, we fix an $M\in\mathbb{Z}$ and consider the graded unital spectral triple $\big{(}\mathcal{O}(\mathbb{C}P_{q}^{r}),L^{2}(\Omega_{M},h),D_{q}\big{)}$ described in Subsection 10.1.

### v2

@@ 10. The quantum spheres and the quantum projective spaces
We are also interested in the $q$-deformed version of the special unitary group $C(SU_{q}(r+1))$ which we refer to as the $C^{*}$-algebraic quantum special unitary group. This compact quantum group was introduced by Stanisław Woronowicz in [Wor87b, Wor87a, Wor88] and it contains the $C^{*}$-algebraic quantum sphere $C(S_{q}^{2r+1})$ as a unital $C^{*}$-subalgebra. Using the description of $C(SU_{q}(r+1))$ from [KlSc97, Chapter 9.2] the relevant map is given by $z_{i}\mapsto u_{r+1,i}$ for all $i\in\{1,\ldots,r+1\}$. The injectivity of this unital *-homomorphism is described in [VaSo90, She97, HoSz02] and a detailed argument can also be found in the review [D’A24, Section 3.3] following the scheme of [MiKa23].
@@ 10.2. Noncommutative metrics on the quantum spheres
We are now going to show that Assumption 7.5 is satisfied in the context of the higher Vaksman-Soibelman quantum spheres. This leads us straight to the last main result of this paper where we show that each of the quantum spheres can be endowed with the structure of a compact quantum metric space arising from noncommutative geometric data. In particular, we generalize one of the main results from [KaKy25] regarding quantum metrics on quantum $SU(2)$. More precisely, our methods yield a new proof of [KaKy25, Theorem 5.23] in the special case where the parameter $t$ (appearing in the statement of [KaKy25, Theorem 5.23]) is equal to one.
It is relevant to clarify that our approach to the spectral geometry of the higher quantum spheres is different from the approach in [ChPa08] for a multitude of reasons. First of all, the approach in [ChPa08] does not have much to do with the pairing between the quantized enveloping algebra and the coordinate algebra for quantum $SU(r+1)$ in the sense that it seems difficult to describe the selfadjoint unbounded operators analyzed in [ChPa08] in terms of this pairing. Another important difference is that in [ChPa08] the authors construct equivariant spectral triples over the higher quantum spheres. As explained in the introduction, our faithfulness to the underlying $q$-geometry, as witnessed by the above mentioned pairing between the quantized enveloping algebra and the coordinate algebra for quantum $SU(r+1)$, prevents us from obtaining unital spectral triples for the higher quantum spheres. Instead, we are forced to work with a different kind of spectral geometric data resulting in two (twisted) *-derivations instead of a single one. These two (twisted) *-derivations are in some sense incompatible because they are twisted with two different twists (where one of the twists is in fact trivial). This has to do with the way the corresponding selfadjoint unbounded operators interact with the coordinate algebra for the quantum sphere. The non-twisted *-derivation is related to the vertical geometric data as witnessed by the circle action and the twisted *-derivation is related to the horizontal geometric data stemming from the unital spectral triple over quantum projective space as reviewed in Theorem 10.2.
Let us provide the relevant details. Our starting point is the $C^{*}$-algebraic quantum sphere $A:=C(S_{q}^{2r+1})$ which is equipped with the strongly continuous circle action $\sigma$ satisfying that $\sigma_{\lambda}(z_{i})=\lambda z_{i}$ for all $i\in\{1,\ldots,r+1\}$. We are moreover considering the coordinate algebra for the quantum sphere, $\mathcal{A}:=\mathcal{O}(S_{q}^{2r+1})$, which is a norm-dense unital *-subalgebra of $C(S_{q}^{2r+1})$. On top of this data, we fix an $M\in\mathbb{Z}$ and consider the graded unital spectral triple $\big{(}\mathcal{O}(\mathbb{C}P_{q}^{r}),L^{2}(\Omega_{M},h),D_{q}\big{)}$ described in Subsection 10.1.